The current price of JCN stock is trading on the NYSE at 131.91. You have figured out that the price goes up 1.33 points on average 55% of the time and goes down .8 points on average 45% of the time. Find the arbitrage of the stock if the time is three months and the rate is 5%, (F=S0*e^(rt), find for F)
Answer #1

The current price of JCN stock is trading on the NYSE at 131.91.

The arbitrage of the stock is = 131.91*e^(0.15) = 131.91 * 1.1618 = 153.253